Out of 20 million people in the population, 9 million people are in the labour force and 3 million are unemployed, therefore the unemployment rate is?
1
Expert's answer
2020-10-08T12:22:49-0400

Unemployment rate = "\frac{number of unemployed people}{total labour force} \times 100"


="\frac{3}{9}\times100"


Unemployment rate= "33\frac{1}{3}%" %